# On the Baltimore Light RailLink into the quantum future

### v1

@@ I Introduction
Only rather recently, employing NISQ for railway scheduling has been demonstrated [9]. In this work, re-scheduling was expressed as a quadratic unconstrained binary optimization (QUBO) problem, which was then solved using a D-Wave quantum annealer. Importantly, while ideally the output of a quantum annealer is the exact optimal solution (corresponding to the ground state of the Ising model used to encode the QUBO), higher-lying excited states are often returned instead. In terms of the original scheduling problem, these higher-lying states may correspond to rail traffic with additional disturbances. See also Ref. [10] for a complementary approach.
In the present work, we demonstrate the utility of NISQ devices by levering the inherent noise present in all such devices as a tool to model stochastic behavior in real-world problems, an idea that will prove especially relevant to the development of hybrid quantum/classical solution methods. Such hybrid methods can tackle significantly larger problems then ether classical or NISQ approach by themselves [11].
Technically, our approach follows the ideas presented in Ref. [12] but with a focus on the practical application of railway re-scheduling. We focus on rail systems with some stochastic features as can be found in tramways, which partially share infrastructure with road traffic. Some research has been performed on this problem [13] and on stochastic optimization of tramway scheduling in general, e.g., in the Hong Kong tram network [14], however literature on this topic is sparse [8] and generally with a different focus such as infrastructure repair scheduling [15] or vehicle and crew rescheduling [16].
In our work, we focus on a specific real-world transportation system: the Baltimore Light RailLink in Baltimore, Maryland  [^1] https://www.mta.maryland.gov/schedule/lightrail visited 13.12.2023. The Baltimore Light RailLink system is an important factor in the economic development in the urban area surrounding the network [18], and from current data on its performance  [^2] https://www.mta.maryland.gov/performance-improvement visited 13.12.2023 indicates that there is a margin for improvement. These properties of the Baltimore Light RailLink infrastructure are in no way unique and are found in other tramway infrastructures (for example that in Karlsruhe, Germany [20]), therefore the conclusions we draw in this work are general.
@@
This article is organized as follows: In Sec. II we present the model of the Baltimore Light RailLink and solution methods for both the quantum annealing and gate-based quantum computing paradigms. In Sec. III we present results from experiments with each of these approaches, using machines from D-Wave and IonQ, alongside simulations. We discuss these results in light of real-life data measured from railway traffic, before presenting our conclusions in Sec. IV. We include several appendices with additional detail, on encoding constraints in QUBO problems (App. A), on a proposed hybrid algorithm that can be applied to larger railway scheduling problems (App. B), on hardware details and embeddings or circuits used for our experiments with a D-Wave quantum annealer (App. C) and IonQ’s gate-based trapped ion computer (App. D), and finally on some tests using IBM’s superconducting gate-based computers (App. E).
@@ II.1.1 Variables
Importantly, the maxmal delay parameter $d_{\max}$ determines the problem size by constraining the range of time variables. We have at most #$\#S\#J$ such time variables, as we do not consider recirculation [27].
@@ II.2 QUBO approach for quantum computing
to determine the minimum objective value $E_{0}=\min_{\vec{x}}E(\vec{x})$ and optimal solution $\vec{x}_{0}=\textrm{argmin}E(\vec{x})$. $Q$ is the matrix encoding the constraints and objective function, which can be taken to be symmetric. The details on encoding the minimal passing time [cf. Eq. (7)] minimal headway [cf. Eq. (5)] and rolling stock circulation [cf. Eq. (6)] constraints as discussed in Section II.1.2 in terms of a quadratic form on the binary decision variables are presented in Appendix A. Ultimately, we expect both the number of qubits ($N$) and the number of non-zero entries in the $Q$ matrix to be proportional to #$\#J\#S$. Hence, for constant $d_{\text{max}}$, the mean degree of the problem graph is expected to be independent on #$\#J$ and #$\#S$.
@@
Once expressed in this form, it is straightforward to input the constraints into the $Q$ matrix of Eq. (11). Following [9], we chose one value $p_{\text{pair}}$ for all quadratic terms in Eq. (13) associated with violations of the constraints, but the generalization is straightforward. The number of these constraints is expected to be linear in #$\#S$$\#J$, as discussed in Section II.1.2 and Appendix A.
@@ III Computational results
From a real Baltimore Light RailLink timetable  [^3] MTA schedule, accessed on 01.12.2023, https://www.mta.maryland.gov/schedule/lightrail we can derive a set of model parameters. For the whole network, these are the minimum headway time $\delta^{\text{headway}}=2$, the minimum preparation time $\delta^{\text{preparation}}=3$, and minimum time in station $\delta^{\text{station}}=1$. For the passage between stations MR and PS, the minimum passing times in the two directions are $\delta^{\text{pass}}_{\text{MR, PS}}=14$ and $\delta^{\text{pass}}_{\text{PS, MR}}=2$.
To model train traffic on this segment, we start with the real Baltimore Light RailLink timetable without any disturbances. Note that in our model we also include several trains whose route either starts from or ends at station PS, though these trains do not run at the current time as station PS is closed due for reconstruction  [^4] MDOT MTA Transit Information, accessed on 01.12.2023, https://www.mta.maryland.gov/schedule/stops/lightrail. This allows us to model rather dense railway traffic, and so we may limit the maximum secondary delay $d_{\rm max}$ which ensures the QUBO representations are not too large.
@@
In total, we consider $30$ distinct scheduling problems. When translated into the QUBO formulation, the smallest problem we consider (with a single train serving $2$ stations and a maximum delay of $2$ minutes) requires $6$ variables to encode. Mapped to a quantum Ising model for quantum annealing, this corresponds to $6$ logical qubits. The largest problem (with $12$ trains, each serving $2$ or $3$ stations, and a maximum delay of $6$ minutes) requires $196$ variables, which is comparable in size to the largest railway scheduling problems solved previously on a D-Wave machine in [9]. Note that since the D-Wave machine does not support arbitrary connectivity between its qubits, there is some overhead in embedding the QUBOs corresponding to our scheduling problems. Figure 2a shows the relationship between the number of logical qubits required for each QUBO we consider and the number of physical qubits necessary after embedding, which can be seen to be approximately linear. Also shown in Fig. 2b is an extrapolation of this scaling to a large enough number of logical qubits to handle problematic real-world railway scheduling problems [7], which are around 60×60$\times$60 × larger than the largest problem we consider in this work (in terms of #$\#J\#S$, e.g. including around 13 stations and 150 trains).
@@ III.1 Quantum annealer
For the exploration of quantum annealing in this work, we have employed the D-Wave Advantage_system6.3 machine with 5614 qubits and a configurable annealing time $T$ (for hardware specifics, see Appendix C). For each scheduling problem, we executed $25000$ runs to build a large distribution of solutions. We did this for each of the $30$ scheduling problems we defined, each run twice with annealing times of $10{\mu s}$ and $1000{\mu s}$ to study the effects on performance.
@@ III.2 Gate-based computers
For the gate-based approach, we solved our scheduling problems with QAOA running on the trapped ion hardware and noisy simulator provided by IonQ. IonQ’s “Aria-1” device provides 25 qubits with all-to-all connectivity, meaning that there is no embedding overhead when solving any given QUBO unlike with D-Wave’s quantum annealers. We solved four non-disturbed scheduling problems with variable counts of 6, 10, 14, and 18 on the Aria-1 device and one disturbed scheduling problem with 18 variables. These problems required the execution of quantum circuits operating on 6 to 18 qubits. For details on these circuits and on the hardware properties, see Appendix D. Taking the largest 18 variable QUBO, the corresponding QAOA circuit with a single-layer ansatz contained 54 single-qubit gates and 30 two-qubit gates. In terms of circuit complexity, the most challenging QUBO to solve is actually the 14 variable instance, which has a more complicated set of constraints leading to the single-layer ansatz compiling to a circuit with 42 single-qubit gates and 63 two-qubit gates.
On the Aria-1 device, the average gate error rates can vary considerably over time, with single-qubit gate errors generally averaging between $0.03\%$ and $0.06\%$ and with two-qubit gate errors generally averaging between $2.1\%$ and $8.6\%$. Further details on the characterization of the device are given in Appendix D. To understand the impact of gate errors on the device performance for our scheduling problems, we solved each QUBO problem using QAOA with both a one- and two-layer ansatz. In principle, the more complicated two-layer variational ansatz should allow the prepared state to better approximate the true ground state of the problem Hamiltonian and therefore should lead to better optimization performance. In practice, however, this theoretical improvement is offset by the increased circuit depth and so increased susceptibility to gate errors and decoherence. Here, for the largest 18 variable QUBO the two-layer QAOA circuit required 90 single- and 60 two-qubit gates. This is approximately double the number of the simpler circuit for the single-layer ansatz.
@@
Figure 6 shows the results of the experiments and simulations solving the largest $18$ variable problem with a two-layer ansatz running on the Aria-1 device. The results are somewhat mixed, and we can not conclude which approach is better in this case.
@@
For one specific two train scheduling problem with a disturbance requiring 18 variables, a comparison of the results obtained with the IonQ Aria-1 device with both ansatze and the results obtained with the D-Wave annealer is presented in Fig. 7 (for an additional point of comparison, we present some results from simulations of IBM’s superconducting devices in Appendix E). In this case, it is clear that the single-layer QAOA ansatz performs better than the two-layer ansatz, especially when the feasible and infeasible spectra overlap. An interesting qualitative observation is that while there is some similarity in the shapes of the histograms of result energies obtained with the different platforms, they are nonetheless visibly different. This is true for the different variational ansatze on the Aria-1 device as well, especially in the distribution of feasible solutions.
Figure 8 continues the comparison between the results obtained with the D-Wave machine and the IonQ simulator running QAOA with the one-layer ansatz, now focusing on the properties of the solutions to the two train 18 variable railway scheduling problem, both with and without disturbances. Here, we show a comparison of the objective values (the tardiness) associated with the feasible solutions returned by each device, alongside the passing times from ${\rm MR}\rightarrow{\rm CS}$ encoded into each solution. We observe that the distribution of passing times produced by the two approaches are very similar, despite the distributions of objective values being quite different.
@@ III.3 Scaling
An important consideration for any proposed application of NISQ devices is scaling behavior, as that will indicate the size of problem instances that can be handled at present and inform what problems are likely to be accessible in the near future. To this end, in Fig. 9 we plot the fraction of schedules returned as solutions to our railway scheduling problems which were feasible (i.e., those which fulfill all constraints listed in Sec. II.2) as a function of the number of physical qubits required for their solution. With both IonQ’s gate-based simulator and D-Wave’s quantum annealer, we unsurprisingly observe that larger problems are more likely to yield purported solutions which violated one or more constraints. For the D-Wave device, we have sufficient statistics to infer that the fraction of feasible solutions decreases exponentially with the number of physical qubits necessary to embed the problem. Naively extrapolating to the sizes necessary to solve problems of practical significance in railway scheduling (e.g., scheduling a crowded metro [7] which is roughly 60×60$\times$60 × larger than the largest problem solved in this work) indicates that such problems will not be accessible even to large quantum devices without some other improvements. Nevertheless, there exist approaches such as that of multilevel combinatorial optimization [22] based on hybrid computation which can allow the solution of problems an order of magnitude larger than current quantum devices. To this end, we examine the statistical distribution of the results produced by our experiments with a goal of using quantum computing as one component of a hybrid algorithm for railway scheduling.

### v2

@@ I Introduction
Only rather recently, employing NISQ for railway scheduling has been demonstrated [9]. In this work, re-scheduling was expressed as a quadratic unconstrained binary optimization (QUBO) problem, which was then solved using a D-Wave quantum annealer – for the conceptual extension of the model to a higher order unconstrained binary optimization (HOBO) approach see [10]. Importantly, while ideally the output of a quantum annealer is the exact optimal solution (corresponding to the ground state of the Ising model used to encode the QUBO), higher-lying excited states are often returned instead. In terms of the original scheduling problem, these higher-lying states may correspond to rail traffic with additional disturbances. Referring to railways see Ref. [11] (high-speed trains timetable optimization on coherent Ising machine quantum simulator), Ref. [12] (rolling stock planning on the quantum annealer for German Railways Network), and Refs. [13, 14] (finding the rail transportation plan on the IBM 32-qubit QASM simulator using QUBO/HOBO and quantum approximate optimization algorithm - QAOA - encoding) for a complementary approaches. However, these studies did not tie the noise of NISQ device with the stochastic behavior of a particular public transportation system.
In the present work, we demonstrate the utility of NISQ devices (and systematically compare annealers and gate-based computers) for the problem of railway scheduling by leveraging the inherent noise present in all such devices as a tool to model stochastic behavior in real-world problems, an idea that will prove especially relevant to the development of hybrid quantum/classical solution methods – see Supplementary Information for pseudo code. Such hybrid methods can tackle significantly larger problems than either the classical or NISQ approach by themselves [15].
Technically, our approach follows the ideas presented in Ref. [16] but with a focus on the practical application of railway re-scheduling. We focus on rail systems with some stochastic features as can be found in tramways, which partially share infrastructure with road traffic. Some research has been performed on this problem [17] and on stochastic optimization of tramway scheduling in general, e.g., in the Hong Kong tram network [18], however literature on this topic is sparse [8] and generally with a different focus such as infrastructure repair scheduling [19] or vehicle and crew rescheduling [20]. In our previous research, non-optimal but practically sound solutions of railway scheduling [21] and analogous Automated Guided Vehicles scheduling [22] problems have been obtained from D-Wave’s proprietary hybrid solvers. Unfortunately, these solvers are closed black boxes and it isn’t easy to assess the actual effort done by the QPU during optimization. Given such experience, in this work we concentrate on pure quantum computing with trackable performance.
In our work, we focus on a specific real-world transportation system: the Baltimore Light RailLink in Baltimore, Maryland [23] The Baltimore Light RailLink system is an important factor in the economic development in the urban area surrounding the network [24], and current data on its performance [25] indicates that there is a margin for improvement. These properties of the Baltimore Light RailLink infrastructure are in no way unique and are found in other tramway infrastructures (for example that in Karlsruhe, Germany [26]), therefore the conclusions we draw in this work are general.
@@
This article is organized as follows: In Sec. II we present the model of the Baltimore Light RailLink and solution methods for both the quantum annealing and gate-based quantum computing paradigms. In Sec. III we present results from experiments with each of these approaches, using machines from D-Wave and IonQ, alongside simulations. We discuss these results in light of real-life data measured from railway traffic, before presenting our conclusions in Sec. IV. Additional details on encoding constraints in QUBO problems, on a proposed hybrid algorithm that can be applied to larger railway scheduling problems, on hardware details and embeddings or circuits used for our experiments with a D-Wave quantum annealer and IonQ’s gate-based trapped ion computer, and on some tests using IBM’s superconducting gate-based computers are available as supplementary material.
@@ II.1.1 Variables
Importantly, the maximal delay parameter $d_{\max}$ determines the problem size by constraining the range of time variables. We have at most #$\#S\#J$ such time variables, as we do not consider recirculation [32].
@@ II.2 QUBO approach for quantum computing
to determine the minimum objective value $E_{0}=\min_{\vec{x}}E(\vec{x})$ and optimal solution $\vec{x}_{0}=\textrm{argmin}E(\vec{x})$. $Q$ is the matrix encoding the constraints and objective function, which can be taken to be symmetric. The details on encoding the minimal passing time [cf. Eq. (7)] minimal headway [cf. Eq. (5)] and rolling stock circulation [cf. Eq. (6)] constraints as discussed in Section II.1.2 in terms of a quadratic form on the binary decision variables are presented in the Supplementary Information. Ultimately, we expect both the number of qubits ($N$) and the number of non-zero entries in the $Q$ matrix to be proportional to #$\#J\#S$. Hence, for constant $d_{\text{max}}$, the mean degree of the problem graph is expected to be independent on #$\#J$ and #$\#S$.
@@
Once expressed in this form, it is straightforward to input the constraints into the $Q$ matrix of Eq. (11). Following [9], we chose one value $p_{\text{pair}}$ for all quadratic terms in Eq. (13) associated with violations of the constraints, but the generalization is straightforward. The number of these constraints is expected to be linear in #$\#S$$\#J$, as discussed in Section II.1.2 and the Supplementary Information.
@@
Recall that if we explicitly require that trains do not meet and overtake at or between stations (see Section II.1.2), a HOBO encoding such as in Ref.[10] would be required. However, such an approach would substantially complicate the problem for the annealer, leading to rather limited practical benefit in our case. For example, in Ref.[10] HOBO problems corresponding to at most $3$ trains and $2$ stations were handled successfully by pure quantum annealing on the D-Wave machine. Nevertheless, we acknowledge, that the HOBO encoding can be more competitive when applying the gate-based quantum computing via the QAOA approach, which opens a particular avenue for further research.
@@ II.4 Model for Gate Quantum Computing
As a final note, our decision to employ QAOA as our algorithm of choice for solving our problem instances using gate-based quantum hardware is motivated primarily by simplicity, efficacy, and ease of comparison to the results of quantum annealing. One of the major limitations of this approach is the limited qubit count available with presently-available gate-based quantum computers. This limitation can be mitigated in several ways, for example by considering hybrid quantum-classical optimizers that only solve subproblems on the quantum hardware (as we discuss in the Supplementary Information), or by considering clever encoding schemes that allow the solution of large problems with few qubits [45, 46, 47]. Such approaches are especially important on current NISQ hardware, as smaller circuits on fewer qubits promise higher accuracy and better optimization performance. An performance advantage due to encoding may also be possible if instead of distilling the scheduling problem to a QUBO, it is encoded as a higher-order binary optimization problem (HOBO). QAOA applied to higher-order problems can require fewer qubits, at the cost of increased circuit depth [48, 49].
Embracing the gate-based paradigm completely, there also exist approaches capable of solving scheduling problems without initially reducing the problem to a QUBO, for instance through quantum versions of the branch-and-bound algorithm [50, 51]. These algorithms pose a challenge for current devices, but in the future may be a promising avenue towards solving large-scale problems with gate-based quantum hardware.
@@ III Computational results
From a real Baltimore Light RailLink timetable [52] we can derive a set of model parameters. For the whole network, these are the minimum headway time $\delta^{\text{headway}}=2$, the minimum preparation time $\delta^{\text{preparation}}=3$, and minimum time in station $\delta^{\text{station}}=1$. For the passage between stations MR and PS, the minimum passing times in the two directions are $\delta^{\text{pass}}_{\text{MR, PS}}=14$ and $\delta^{\text{pass}}_{\text{PS, MR}}=2$.
To model train traffic on this segment, we start with the real Baltimore Light RailLink timetable without any disturbances. Note that in our model we also include several trains whose route either starts from or ends at station PS, though these trains do not run at the current time as station PS is closed due for reconstruction [53]. This allows us to model rather dense railway traffic, and so we may limit the maximum secondary delay $d_{\rm max}$ which ensures the QUBO representations are not too large.
@@
In total, we consider $30$ distinct scheduling problems. When translated into the QUBO formulation, the smallest problem we consider (with a single train serving $2$ stations and a maximum delay of $2$ minutes, where the delay is counted from zero) requires $6$ variables to encode. Mapped to a quantum Ising model for quantum annealing, this corresponds to $6$ logical qubits. The largest problem (with $12$ trains, each serving $2$ or $3$ stations, and a maximum delay of $6$ minutes) requires $196$ variables, which is comparable in size to the largest railway scheduling problems solved previously on a D-Wave machine in [9]. Note that since the D-Wave machine does not support arbitrary connectivity between its qubits, there is some overhead in embedding the QUBOs corresponding to our scheduling problems. For a deeper discussion, see the Supplementary Information. Fig. 2a shows the relationship between the number of logical qubits required for each QUBO we consider and the number of physical qubits necessary after embedding, which can be seen to be approximately linear. Also shown in Fig. 2b is an extrapolation of this scaling to a large enough number of logical qubits to handle problematic real-world railway scheduling problems [7], which are around 60×60$\times$60 × larger than the largest problem we consider in this work (in terms of #$\#J\#S$, e.g. including around 13 stations and 150 trains).
@@ III.1 Quantum annealer
For the exploration of quantum annealing in this work, we have employed the D-Wave Advantage_system6.3 machine with 5614 qubits and a configurable annealing time $T$ (for hardware specifics, see the Supplementary Information). For each scheduling problem, we executed $25000$ runs to build a large distribution of solutions. We did this for each of the $30$ scheduling problems we defined, each run twice with annealing times of $10{\mu s}$ and $1000{\mu s}$ to study the effects on performance.
@@ III.2 Gate-based computers
For the gate-based approach, we solved our scheduling problems with QAOA running on the trapped ion hardware and noisy simulator provided by IonQ. IonQ’s “Aria-1” device provides 25 qubits with all-to-all connectivity, meaning that there is no embedding overhead when solving any given QUBO unlike with D-Wave’s quantum annealers. We solved four non-disturbed scheduling problems with variable counts of 6, 10, 14, and 18 on the Aria-1 device and one disturbed scheduling problem with 18 variables. These problems required the execution of quantum circuits operating on 6 to 18 qubits. For details on these circuits and on the hardware properties, see the Supplementary Information. Taking the largest 18 variable QUBO, the corresponding QAOA circuit with a single-layer ansatz contained 54 single-qubit gates and 30 two-qubit gates. In terms of circuit complexity, the most challenging QUBO to solve is actually the 14 variable instance, which has a more complicated set of constraints leading to the single-layer ansatz compiling to a circuit with 42 single-qubit gates and 63 two-qubit gates.
On the Aria-1 device, the average gate error rates can vary considerably over time, with single-qubit gate errors generally averaging between $0.03\%$ and $0.06\%$ and with two-qubit gate errors generally averaging between $2.1\%$ and $8.6\%$. Further details on the characterization of the device are given in the Supplementary Information. To understand the impact of gate errors on the device performance for our scheduling problems, we solved each QUBO problem using QAOA with both a one- and two-layer ansatz. In principle, the more complicated two-layer variational ansatz should allow the prepared state to better approximate the true ground state of the problem Hamiltonian and therefore should lead to better optimization performance. In practice, however, this theoretical improvement is offset by the increased circuit depth and so increased susceptibility to gate errors and decoherence. Here, for the largest 18 variable QUBO the two-layer QAOA circuit required 90 single- and 60 two-qubit gates. This is approximately double the number of the simpler circuit for the single-layer ansatz.
@@
Fig. 6 shows the results of the experiments and simulations solving the largest $18$ variable problem with a two-layer ansatz running on the Aria-1 device. The results are somewhat mixed, and we can not conclude which approach is better in this case.
@@
For one specific two train scheduling problem with a disturbance requiring 18 variables, a comparison of the results obtained with the IonQ Aria-1 device with both ansatze and the results obtained with the D-Wave annealer is presented in Fig. 7 (for an additional point of comparison, we present some results from simulations of IBM’s superconducting devices in the Supplementary Information). In this case, it is clear that the single-layer QAOA ansatz performs better than the two-layer ansatz, especially when the feasible and infeasible spectra overlap. An interesting qualitative observation is that while there is some similarity in the shapes of the histograms of result energies obtained with the different platforms, they are nonetheless visibly different. This is true for the different variational ansatze on the Aria-1 device as well, especially in the distribution of feasible solutions. To explain the potentially better performance in the cases with overlapping spectra, recall that high penalty values can reshape the low energy part of the optimization landscape. This potentially introduces gaps before good feasible solutions, and hence may trap the quantum optimization in higher energy states – see the Supplementary Information for further discussion including the entire spectra (split and overlapping) of selected QUBO problems.
Fig. 8 continues the comparison between the results obtained with the D-Wave machine and the IonQ simulator running QAOA with the one-layer ansatz, now focusing on the properties of the solutions to the two train 18 variable railway scheduling problem, both with and without disturbances. Here, we show a comparison of the objective values (the tardiness) associated with the feasible solutions returned by each device, alongside the passing times from ${\rm MR}\rightarrow{\rm CS}$ encoded into each solution. We observe that the distribution of passing times produced by the two approaches are very similar, despite the distributions of objective values being quite different.
@@ III.3 Scaling
An important consideration for any proposed application of NISQ devices is scaling behavior, as that will indicate the size of problem instances that can be handled at present and inform what problems are likely to be accessible in the near future. To this end, in Fig. 9 we plot the fraction of schedules returned as solutions to our railway scheduling problems which were feasible (i.e., those which fulfill all constraints listed in Sec. II.2) as a function of the number of physical qubits required for their solution. With both IonQ’s gate-based simulator and D-Wave’s quantum annealer, we unsurprisingly observe that larger problems are more likely to yield purported solutions which violated one or more constraints. For the D-Wave device, we have sufficient statistics to infer that the fraction of feasible solutions decreases exponentially with the number of physical qubits necessary to embed the problem. Bear in mind that the translation of the scheduling problem to QUBO significantly increases the number of variables due to the time discretization (Eq. (10)), which is not necessary in the original ILP formulation. To reflect this in practice, we have solved the deterministic ILP problem with CPLEX (Python API version: 22.1.1.0 on the local CPU). In our case, the CPLEX solution was always optimal and the CPLEX computational time was short ($0.0014$ s to $0.08$ s), pinpointing the better performance of the ILP approach. However, our goal was to asses the accessibility of various NISQ devices to model actual railway stochastic system rather than to compare runtimes in finding optimal solutions.
Naively extrapolating to the sizes necessary to solve problems of practical significance in railway scheduling (e.g., scheduling a crowded metro [7] which is roughly 60×60$\times$60 × larger than the largest problem solved in this work) indicates that such problems will not be accessible even to large quantum devices without some other improvements. Nevertheless, there exist approaches such as that of multilevel combinatorial optimization [27] based on hybrid computation which can allow the solution of problems an order of magnitude larger than current quantum devices. To this end, we examine the statistical distribution of the results produced by our experiments with a goal of using quantum computing as one component of a hybrid algorithm for railway scheduling.